Twelve more than the square of the difference of four times a number and eight in algebraic expression.
Let x represent the unknown number.

Found 2 solutions by JoelSchwartz, ewatrrr:
Answer by JoelSchwartz(130)

12+(4x-8)^2
12+(4x-8)(4x-8)
12+16x^2-32x-32x+64
16x^2-64x+76

Answer by ewatrrr(24785)

(4n -8)^2 + 12 = 16n^2 - 64n + 64 + 12 = 16n^2 - 64n + 76 = 4(4n^2 - 16n + 19)